OpenStudy (anonymous):

Which rule will find the nth term of the arithmetic sequence -58, -65, -72, -86, …? an = -65 − 7n an = -65 + 7n an = 51 − 7n an = -51 + 7n an = -51 − 7n

OpenStudy (gorv):

do u see any pattern in these term??

OpenStudy (gorv):

a number is subtracted or added to previous term??

OpenStudy (gorv):

hint it is an arithmatic progression (A.P)

OpenStudy (anonymous):

i still dont understand though @gorv

OpenStudy (gorv):

u studied arithmatic progression??

OpenStudy (anonymous):

yes in the beggining of the year

OpenStudy (gorv):

can u tell me the common difference ??

OpenStudy (gorv):

@brittany.hodges

OpenStudy (anonymous):

there all negitive @gorv

OpenStudy (gorv):

just apply what u do for common difference like two number -3 and -5 difference = -5-(-3)=-5+3=-2 do like this

OpenStudy (anonymous):

thanks i got it @gorv

OpenStudy (gorv):

welcome :)